OpenStudy (trojanpoem):

if A,B are two matrices that verify the relationship AB = BA -> (1) prove that: (AB)^n = A^nB^n I tried: (AB)^n = (AB)* (AB) * (AB) * (AB) * (AB) .... (AB) from -> (1) AB * BA * AB * BA *AB * BA.... BA AB^2A^2B^2A^2B^2....BA from 1 AB^2 * B^2A^2 * A^2B^ ... B = A^nB^n Is there a math-friendly way of writing that ? or a better solution ?

OpenStudy (dan815):

http://prntscr.com/8r59a2

OpenStudy (empty):

I don't know if this question is really worth spending too much time on it, but hey it does seem like there's gotta be a better way. So how about like this, if I show it's true for one case then we can see it'll be true for all cases: \[A^nB^nAB=A^nAB^nB=A^{n+1}B^{n+1}\] So I think since \(ABAB=A^2B^2\) as our base case then for all higher cases we can always step up to it with the thing I said.

OpenStudy (dan815):

well AB = BA so xD

OpenStudy (trojanpoem):

@Empty , My professor denies using induction.

OpenStudy (dan815):

so its kinda pointless

OpenStudy (dan815):

AB AB AB AB AB A....... BA BA BA BA BA ....B

OpenStudy (empty):

Eh whatever then you're not gonna get anything better I think lol

OpenStudy (trojanpoem):

@dan815 , That's what I did, but I doubt it's understood-able . I am looking for a notation or a better solution

OpenStudy (dan815):

the first post u mean?

OpenStudy (dan815):

thsi is the same thing as kais... u can keep switchinb AB to BA and keep getting more and more A^k and B^k

OpenStudy (trojanpoem):

I posted a solution associated with the question itself.

OpenStudy (trojanpoem):

Yea, I understand the idea. But will I just write AB * BA * AB * BA... BA = A^n B^n

OpenStudy (dan815):

no u wont do that

OpenStudy (dan815):

you have to show a strategy how to get more A^ks on the left and B^ks on the right

OpenStudy (trojanpoem):

That's more readable, but why n+1 ?

OpenStudy (dan815):

A (BA ,,BA BA BA)^n B =A^1 (AB AB AB... AB)^n B^1 A (AB AB .... AB AB) B = A^2 (BA .... BA BA)^n-1 B^2

OpenStudy (dan815):

u lose an A and B from the middle part every step

OpenStudy (dan815):

this can be your induction step

OpenStudy (empty):

lol you shouldn't have said that

OpenStudy (trojanpoem):

Said what ?

OpenStudy (trojanpoem):

Do you mean the induction word xD ?

OpenStudy (dan815):

show A^k (AB)^(n+1) B^k = A^(k+1) (AB)^n B^(k+1)

OpenStudy (dan815):

i think its fine your teacher shouldnt have problems with this

OpenStudy (trojanpoem):

Professor not teacher, He is enthusiast about giving zeros out.

OpenStudy (dan815):

well if u want to be a little more correct

OpenStudy (dan815):

u can write this expression so that there is a relationship between k and n

OpenStudy (dan815):

A^k (AB)^(n-k) B^k = A^(k+1) (AB)^(n-k-1) B^(k-1)

OpenStudy (dan815):

this way u can show one induction step and say that when k reaches n you will be left with A^n B^n

OpenStudy (trojanpoem):

Well, I wish he'd like this, I hate zeros xD

OpenStudy (dan815):

just write this out formally

OpenStudy (trojanpoem):

Show me the formal way of writing them.

OpenStudy (dan815):

xD

OpenStudy (trojanpoem):

it*

OpenStudy (dan815):

that too much work

OpenStudy (dan815):

u get the idea, just do the best u can